Show that U(8)/is isomorphic to U(12)


1
Expert's answer
2021-08-10T14:24:10-0400

U(8): 1, 3, 5, 7 , "3^2=5^2=7^2" =1 mod 8


U(12): 1, 5, 7, 11, "5^2=7^2=11^2" =1 mod 12


The map "\sigma:U(8)->U(12): \sigma(3)=5, \sigma(5)=7,\sigma(7)=11"

is "1-1," onto and takes the Cayley table of U(8) to that of U(12) so it is isomorphism.